# Quantum Chemical and Trajectory Surface Hopping Molecular Dynamics Study of Iodine-based BODIPY Photosensitizer

### v1

@@ 3.2 Benchmarking TD-DFT functionals
The commonly quoted El-Sayed rules which were originally formulated for diazenes 95, nitrogen heterocyclics 96, carbonyls and halogenated aromatics 97 and later reportedly generalized to (either the same or) some other types of molecules are nothing but a special case of the much more general Slater–Condon rules outlined above and will therefore not be referred to further in this text.
It is evident that the magnitudes of SOC between the lowest-lying excited singlets and triplets, all dominated by $(\pi,\pi^{*})$ excitations, are very likely to increase as soon as the iodinated BODIPY derivative loses planarity—mainly due to the unavoidable mixing of the originally $(n,\pi^{*})$ and $(\pi,\sigma^{*})$ excitations into the ‘zero-order’ $(\pi,\pi^{*})$ states. This notorious phenomenon 98, 99 can be lucidly demonstrated e.g. by plotting the SOC magnitudes in the initial state, typically S1, against the dimensionless (scaled) normal-mode-projected displacement from the planar equilibrium geometry along the normal coordinate of a suitably chosen out-of-plane vibration. This is almost exactly what we did in Fig. 3 for monoiodinated BODIPY in the position 2 (in the ground-state geometry) and its softest normal mode (with harmonic vibrational frequency of about 20–30 cm-1) basically corresponding to bending of the entire molecule along the line connecting the atoms B and C(8), except that we artificially set the same mass to all atoms of the molecule (to effectively restrict the hydrogen atoms in their otherwise rather sweeping harmonic motion) and plotted the C(2)–B–C(8)–C(6) dihedral angle instead of the normal-mode-projected displacement on abscissa.
@@
Accordingly, the existence of a thermally accessible minimum energy crossing point (MEXP) on the crossing seam of the initial singlet and some close-lying triplet located at a non-planar geometry would suggest that the corresponding ISC rate may be enhanced. 37 Obviously, a sufficiently long and accurate direct (on-the-fly) spin-forbidden NAMD simulation would provide the definitive answer. Such a simulation naturally captures not only the geometry dependence of SOC, but also the so-called spin–vibronic coupling (including its own geometry dependence) which is a joint effect of SOC and NAC having its origin in the second order of time-dependent perturbation theory. 100, 101 Actually, this classical terminology is probably going to be forgotten as the term spin–vibronic coupling is now frequently being (mis)used for essentially everything beyond single-point SOC.
@@
In the rest of our study, only B3LYP and BHLYP functionals (the latter exclusively within TDA) are used since analytical TD-DFT gradients for the M06-2X functional are not available in Turbomole v7.0.1. Actually, the use of TDA has certain advantages. Unlike full TD-DFT, TD-DFT(TDA) does not suffer from numerical instabilities near conical intersections, making the method comparatively more stable in these regions. 102, 103 It can thus at least approximately reproduce the topology of the potential energy surface (PES) as obtained by multireference methods. Moreover, TDA offers a remedy for triplet instabilities that impact on TD-DFT. 104
@@ 3.3 Potential energy cuts
To facilitate the comparison, we have also calculated the so-called non-parallelities (NPs) between the potential energy curves. 105. For two curves, $f(R)$ and $g(R)$, NP is defined as
@@ 3.4 Absorption spectra
Simulated absorption spectra of BODIPY and Br-BODIPY from our previous study 38, and of I-BODIPY studied here, all calculated at the TD-DFT/B3LYP/dhf-TZVP level of theory, are displayed in Fig. 4.
@@
It is evident that the spectrum of I-BODIPY shows a remarkable red shift with respect to the spectrum of the unsubstituted BODIPY, quite similar to that of Br-BODIPY, and the spectra of I-BODIPY and Br-BODIPY thus closely overlap. This can be explained by the fact that the (essentially $\pi$ bonding) AO 5p of iodine and AO 4p of bromine, both perpendicular to the plane of the rings, destabilize HOMO and LUMO of the pyrrole radical (unequally, increasing the energy of HOMO slightly more than the energy of LUMO) by their positive mesomeric effect, as we already discussed previously 38. Despite the fact that the mesomeric effect is somewhat stronger for the bromine than for the iodine substituent (mainly due to the smaller covalent radius of the bromine atom), the calculated HOMO/LUMO gap and red shift are virtually the same. The overall stabilization of the frontier MOs, roughly equal for HOMO and LUMO, by the negative inductive effect of the halogen substituents is nevertheless greater for bromine than for iodine due to the higher electronegativity of the former. The shapes and energies of HOMO and LUMO of I-BODIPY are given in Figs. 5 and 6.
Figure 5: HOMO of I-BODIPY, orbital energy -5.98 eV.
Figure 6: LUMO of I-BODIPY, orbital energy -3.07 eV.
@@
The calculated vertical excitation energies and oscillator strengths suggest that I-BODIPY could be an efficient photosensitizer. It has two spectroscopically active excited states, the first and second excited singlets, the optical transitions to which exhibit considerable oscillator strengths regardless of whether they are calculated by ADC(2) or TD-DFT (cf. Tables 1, S2a and S3a). Moreover, the vertical $\Delta E$(T1–S0) energy gap is by a decent margin larger than 0.9 eV, the energy necessary to activate O2 from its triplet ground state ${}^{3}\Sigma_{g}$ to its excited singlet state ${}^{1}\Delta_{g}$106
@@ 3.5 Ultrafast excited-state dynamics
To model the slow spin-forbidden relaxation processes in I-BODIPY, we employed the accelerated NAMD technique 80 as described above. Three sets of MD trajectories were calculated, each set using a different scaling factor of either 2, 3.5 or 5. Each set consisted of 20 trajectories with the initial conditions (nuclear coordinates and momenta) sampled from the Wigner distribution of the quantum harmonic oscillator 107 in the optimized ground-state geometry at 298 K. All trajectories were started in the first excited singlet $\text{S}_{1}$. The quantum chemical method used was TD-DFT(TDA)/BHLYP-D3/dhf-TZVP. The D3 empirical correction for dispersion was applied as dispersion effects may become significant outside the Franck–Condon region 108. We included in the simulations the singlets S0 through S6 and triplets T1 through T10 to cover all the low-lying electronic states that could possibly affect the radiationless deactivation of I-BODIPY in its first excited singlet, see Tables S7a–c. The unusually large number of the embraced higher excited states dominated mainly by $(n,\pi^{*})$, $(\pi,\sigma^{*})$, and $(n,\sigma^{*})$ excitations is due to their extraordinarily strong SOC with the lowest-lying $(\pi,\pi^{*})$ states and/or among themselves, see Table S19.
@@
Strong SOC between $(n,\pi^{*})$ and $(\pi,\pi^{*})$ states in unsaturated organic molecules containing nitrogen, oxygen, or halogen atoms is a notorious phenomenon described by the well-known El-Sayed rules. 95, 96, 97 For organic molecules in which iodine atoms are directly attached to carbon atoms building cyclic $\pi$-electron rings it has already been reported that electronic states with the largest SOC are those which involve transition of an electron into the C–I $\sigma$ antibonding MO either from a $\pi$ bonding MO of the $\pi$-electron ring or from the nonbonding orbital $n$, essentially the AO 5p of the iodine atom, residing on the iodine atom itself 109, 110. Embracing the first six excited singlets and ten triplets in the dynamics will thus get into the game not only the lowest-lying $(\pi,\pi^{*})$ and $(n,\pi^{*})$ states, but also all the other important, mainly $(\pi,\sigma^{*})$ and $(n,\sigma^{*})$, states that can participate in successful ISCs.
@@
The (classical) Newton’s equations of motion for nuclei were integrated with a time step of 0.5 fs by the velocity-Verlet algorithm while the (quantum) time-dependent Schrödinger equation for electrons was integrated with a (20 times shorter) time step of 0.025 fs by a unitary propagator algorithm using linearly interpolated energies and couplings in the substeps. The simulations were caried out in the diagonal representation, i.e., in the spin-adiabatic basis of the eigenfunctions of the two-component psuedorelativistic Hamiltonian computed by quasidegenerate perturbation theory (QDPT) with the spin-orbit Hamiltonian taken for the perturbation operator, using the 3-step integrator approach. 66, 65 The employed first-order QDPT is nothing but a severely limited spin–orbit CI, namely the diagonalization of the matrix of the two-component pseudorelativistic Hamiltonian in the truncated spin-diabatic basis of the eigenfunctions of the (scalar relativistic) spinless Hamiltonian, as a result of which the zero-order QDPT wave functions are just linear combinations of the singlets and (individual components of) triplets included in the basis.
@@
The maximum simulation time was 200 fs for the trajectories calculated using the scaling factors of 2 and 3.5, and 100 fs for the trajectories calculated using the scaling factor of 5. The resulting average populations of singlets and triplets (back-transformed from the spin-adiabatic basis) in the excited-state dynamics of I-BODIPY started in S1 are depicted in Figs. 7, 8 and 9 (for the scaling factors of 2, 3.5 and 5).
@@
The time constant of the overall triplet population has been estimated by fitting scaled and shifted complimentary exponential decay curves to the purple data points in Figs. 7, 8 and 9, yielding $\tau_{2}^{\text{T}}=441.66$ fs, $\tau_{3.5}^{\text{T}}=28.44$ fs, and $\tau_{5}^{\text{T}}=12.44$ fs. Extrapolation of these time constants to $\alpha=1$ gives $\tau_{1}^{\text{T}}=6.06$ ps (as shown in Fig. 10). The inverse of $\tau_{1}^{\text{T}}$ is the ISC rate constant $k_{\text{ISC}}=1.73\times 10^{11}$ s-1.
@@
The lifetime of $\text{S}_{1}$ has been estimated by fitting scaled and shifted exponential decay curves to the yellow data points in Figs. 7, 8 and 9 (see also Figs. S25, S26 and S27), yielding $\tau_{2}^{\text{S}_{1}}=179.1$ fs, $\tau_{3.5}^{\text{S}_{1}}=26.3$ fs, and $\tau_{5}^{\text{S}_{1}}=8.8$ fs. Extrapolation of these lifetimes to $\alpha=1$ gives $\tau_{1}^{\text{S}_{1}}=1.73$ ps (as shown in Fig. S28).
@@
However, it should be noted that the data shown in Figs. S38 and 12 are not completely accurate or even correct. The problem is that by picking out only the time steps at which surface hops between spin-adiabatic states take place we obviously miss the transitions between spin-diabatic states which occur by continuous change of their CI expansion coefficients in the (one and only) active spin-adiabatic state. Actually, sufficiently close to a conical intersection of two spin-adiabatic states (crossing seam of two spin-diabatic states) a hop between spin-adiabatic states may in fact represent staying in a given spin-diabatic state (either singlet or triplet) and vice versa, staying in a given spin-adiabatic state may represent a hop between spin-diabatic states (from singlet to triplet or the other way around). Another undesired consequence of our slightly biased approach are the non-empty histograms on the diagonal of the matrix on Fig. S38.

### v2

@@ 3.2 Benchmarking TD-DFT functionals
The commonly quoted El-Sayed rules which were originally formulated for diazenes 101, nitrogen heterocyclics 102, carbonyls and halogenated aromatics 103 and later reportedly generalized to (either the same or) some other types of molecules are nothing but a special case of the much more general Slater–Condon rules outlined above and will therefore not be referred to further in this text.
It is evident that the magnitudes of SOC between the lowest-lying excited singlets and triplets, all dominated by $(\pi,\pi^{*})$ excitations, are very likely to increase as soon as the iodinated BODIPY derivative loses planarity—mainly due to the unavoidable mixing of the originally $(n,\pi^{*})$ and $(\pi,\sigma^{*})$ excitations into the ‘zero-order’ $(\pi,\pi^{*})$ states. This notorious phenomenon 104, 105 can be lucidly demonstrated e.g. by plotting the SOC magnitudes in the initial state, typically S1, against the dimensionless (scaled) normal-mode-projected displacement from the planar equilibrium geometry along the normal coordinate of a suitably chosen out-of-plane vibration. This is almost exactly what we did in Fig. 12 for monoiodinated BODIPY in the position 2 (in the ground-state geometry) and its softest normal mode (with harmonic vibrational frequency of about 20–30 cm-1) basically corresponding to bending of the entire molecule along the line connecting the atoms B and C(8), except that we artificially set the same mass to all atoms of the molecule (to effectively restrict the hydrogen atoms in their otherwise rather sweeping harmonic motion) and plotted the C(2)–B–C(8)–C(6) dihedral angle instead of the normal-mode-projected displacement on abscissa.
@@
Accordingly, the existence of a thermally accessible minimum energy crossing point (MEXP) on the crossing seam of the initial singlet and some close-lying triplet located at a non-planar geometry would suggest that the corresponding ISC rate may be enhanced. 37 Obviously, a sufficiently long and accurate direct (on-the-fly) spin-forbidden NAMD simulation would provide the definitive answer. Such a simulation naturally captures not only the geometry dependence of SOC, but also the so-called spin–vibronic coupling (including its own geometry dependence) which is a joint effect of SOC and NAC having its origin in the second order of time-dependent perturbation theory. 106, 107 Actually, this classical terminology is probably going to be forgotten as the term spin–vibronic coupling is now frequently being (mis)used for essentially everything beyond single-point SOC.
@@
In the rest of our study, only B3LYP and BHLYP functionals (the latter exclusively within TDA) are used since analytical TD-DFT gradients for the M06-2X functional are not available in Turbomole v7.0.1. Actually, the use of TDA has certain advantages. Unlike full TD-DFT, TD-DFT(TDA) does not suffer from numerical instabilities near conical intersections, making the method comparatively more stable in these regions. 108, 109 It can thus at least approximately reproduce the topology of the potential energy surface (PES) as obtained by multireference methods. Moreover, TDA offers a remedy for triplet instabilities that impact on TD-DFT. 110
@@ 3.3 Potential energy cuts
To facilitate the comparison, we have also calculated the so-called non-parallelities (NPs) between the potential energy curves. 111. For two curves, $f(R)$ and $g(R)$, NP is defined as
@@
Despite the fact that we have eventually selected exchange-correlation functionals more or less suitable for the subsequent NEA absorption spectra and mixed quantum-classical TSH MD simulations on I-BODIPY, the search for a TD-DFT functional ideal for these purposes has evidently not come to its end.
@@ 3.4 Absorption spectra
Simulated absorption spectra of BODIPY and Br-BODIPY from our previous study 38, and of I-BODIPY studied here, all calculated at the TD-DFT/B3LYP/dhf-TZVP level of theory, are displayed in Fig. 13. The simulations have been carried out using the nuclear ensemble approach (NEA), in which the photoabsorption cross sections are averaged over a set of configuration-space points either sampled from an appropriate phase-space distribution or collected in the course of a ground-state MD simulation. The cross section for a given geometry is a continuous function (of photon energy) proportional to the sum over the included excited states (i.e., transitions) of the products of the oscillator strength of each transition and a sharp (usually Gaussian) line-shape function centered at the vertical excitation energy of the transition. 60 Absorption spectra obtained in this way clearly do not show vibrational structure of the bands. Each calculation was based on 1000 nuclear configurations sampled from the Wigner distribution of a canonical ensemble of the noninteracting quantum harmonic oscillators 112 in the optimized ground-state geometry at 298 K while the three lowest-lying excited singlets of I-BODIPY were included.
@@
It is evident that the spectrum of I-BODIPY shows a remarkable red shift with respect to the spectrum of the unsubstituted BODIPY, quite similar to that of Br-BODIPY, and the spectra of I-BODIPY and Br-BODIPY thus closely overlap. This can be explained by the fact that the (essentially $\pi$ bonding) AO 5p of iodine and AO 4p of bromine, both perpendicular to the plane of the rings, destabilize HOMO and LUMO of the pyrrole radical (unequally, increasing the energy of HOMO slightly more than the energy of LUMO) by their positive mesomeric effect, as we already discussed previously 38. Despite the fact that the mesomeric effect is somewhat stronger for the bromine than for the iodine substituent (mainly due to the smaller covalent radius of the bromine atom), the calculated HOMO/LUMO gap and red shift are virtually the same. The overall stabilization of the frontier MOs, roughly equal for HOMO and LUMO, by the negative inductive effect of the halogen substituents is nevertheless greater for bromine than for iodine due to the higher electronegativity of the former. The shapes and energies of HOMO and LUMO of I-BODIPY are given in Figs. 7 and 8.
@@
The calculated vertical excitation energies and oscillator strengths suggest that I-BODIPY could be an efficient photosensitizer. It has two spectroscopically active excited states, the first and second excited singlets, the optical transitions to which exhibit considerable oscillator strengths regardless of whether they are calculated by ADC(2) or TD-DFT (cf. Tables 1, S2a and S3a). Moreover, the vertical $\Delta E$(T1–S0) energy gap is by a decent margin larger than 0.9 eV, the energy necessary to activate O2 from its triplet ground state ${}^{3}\Sigma_{g}$ to its excited singlet state ${}^{1}\Delta_{g}$113
@@ 3.5 Ultrafast excited-state dynamics
To model the slow spin-forbidden relaxation processes in I-BODIPY, we employed the accelerated NAMD technique 86 as described above. Three sets of FSSH MD trajectories were calculated, each set using a different scaling factor of either 2, 3.5 or 5. Each set consisted of 20 trajectories with the initial conditions (nuclear coordinates and momentasampled from the Wigner distribution of a canonical ensemble of the noninteracting quantum harmonic oscillators 112 in the optimized ground-state geometry at 298 K. All trajectories were started in the first excited singlet $\text{S}_{1}$ (or, more precisely, in a quasirelativistic state which had the greatest overlap with the scalar relativistic singlet S1). The quantum chemical method used was TD-DFT(TDA)/BHLYP-D3/dhf-TZVP. The D3 empirical correction for dispersion was applied as dispersion effects may become significant outside the Franck–Condon region 114. We included in the simulations the singlets S0 through S6 and triplets T1 through T10 to cover all the low-lying electronic states that could possibly affect the radiationless deactivation of I-BODIPY in its first excited singlet, see Tables S7a–c. The unusually large number of the embraced higher excited states dominated mainly by $(n,\pi^{*})$, $(\pi,\sigma^{*})$, and $(n,\sigma^{*})$ excitations is due to their extraordinarily strong SOC with the lowest-lying $(\pi,\pi^{*})$ states and/or among themselves, in some cases up to 3 orders of magnitude stronger than the average SOC between the lowest-lying $(\pi,\pi^{*})$ singlets and triplets, see Table S19.
@@
Strong SOC between $(n,\pi^{*})$ and $(\pi,\pi^{*})$ states in unsaturated organic molecules containing nitrogen, oxygen, or halogen atoms is a notorious phenomenon described by the well-known El-Sayed rules. 101, 102, 103 For organic molecules in which iodine atoms are directly attached to carbon atoms building cyclic $\pi$-electron rings it has already been reported that electronic states with the largest SOC are those which involve transition of an electron into the C–I $\sigma$ antibonding MO either from a $\pi$ bonding MO of the $\pi$-electron ring or from the nonbonding orbital $n$, essentially the AO 5p of the iodine atom, residing on the iodine atom itself 115, 116. Embracing the first six excited singlets and ten triplets in the dynamics will thus get into the game not only the lowest-lying $(\pi,\pi^{*})$ and $(n,\pi^{*})$ states, but also all the other important, mainly $(\pi,\sigma^{*})$ and $(n,\sigma^{*})$, states that can participate in successful ISCs.
@@
The (classical) Newton’s equations of motion for nuclei were integrated with a time step of 0.5 fs by the velocity-Verlet algorithm while the (quantum) time-dependent Schrödinger equation for electrons was integrated with a (20 times shorter) time step of 0.025 fs by a unitary propagator algorithm using linearly interpolated energies and couplings in the substeps. The simulations were caried out in the diagonal representation, i.e., in the spin-adiabatic basis of the eigenfunctions of the two-component pseudorelativistic Hamiltonian computed by quasidegenerate perturbation theory (QDPT) with the spin-orbit Hamiltonian taken for the perturbation operator, using the 3-step integrator approach. 71, 70 The employed first-order QDPT is nothing but a severely limited spin–orbit CI, namely the diagonalization of the matrix of the two-component pseudorelativistic Hamiltonian in the truncated spin-diabatic basis of the eigenfunctions of the (scalar relativistic) spinless Hamiltonian, as a result of which the zero-order QDPT wave functions are just linear combinations of the singlets and (individual components of) triplets included in the basis.
Let, at each 0.5 fs time step,
${\bf HU}={\bf UE},$
where H is the complex Hermitian matrix of the two-component pseudorelativistic Hamiltonian $\hat{H}$ in the truncated (and possibly orthonormalized) basis of spin-diabatic states (singlets and individual components of triplets), E is the real diagonal matrix of its eigenvalues (quasirelativistic state energies), and U is the complex unitary matrix of its eigenvectors. Then, obviously,
${\bf U}^{-1}{\bf HU}={\bf U}^{+}\langle{\bf\Theta}|\hat{H}|{\bf\Theta}\rangle{\bf U}=\langle{\bf\Phi}|\hat{H}|{\bf\Phi}\rangle={\bf E}$
while
${\bf\Phi}={\bf\Theta U},$
where $\bf\Theta$ and $\bf\Phi$ are the row vectors of the spin-diabatic and spin-adiabatic bases, respectively.
The FSSH MD simulation provides a sequence of normalized column vectors ${\bf c}(t)$ of the complex expansion coefficients of the sought time-dependent wave function $\Psi(t)$ in the spin-adiabatic basis ${\bf\Phi}(t)$,
$\Psi(t)={\bf\Phi}(t){\bf c}(t)={\bf\Theta}(t){\bf U}(t){\bf c}(t)={\bf\Theta}(t){\bf b}(t),$
where
${\bf b}(t)={\bf U}(t){\bf c}(t)$
must be the normalized column vector of the complex expansion coefficients of $\Psi(t)$ in the spin-diabatic basis ${\bf\Theta}(t)$. The squares $|b_{i}(t)|^{2}\leq 1$ of the absolute values of the elements of ${\bf b}(t)$ are then the singlet and triplet state populations along a certain trajectory, and averaged over the ensemble of trajectories (while summed over the three components of each triplet) they give the final populations shown below.
From the exposition it is clear that not only $\Psi(t)$ and ${\bf\Phi}(t)$, but also ${\bf\Theta}(t)$ evolves in time, and spin-diabatic states of the same spin multiplicity but different characters may thus swap ordinal indices (or just mix) during the simulation, cf. Tables S7a–c. This must always be borne in mind when interpreting the results.
@@
The maximum simulation time was 200 fs for the trajectories calculated using the scaling factors of 2 and 3.5, and 100 fs for the trajectories calculated using the scaling factor of 5. The resulting average populations of singlets and triplets (back-transformed from the spin-adiabatic basis) in the excited-state dynamics of I-BODIPY started in S1 are depicted in Figs. 14, 15 and 16 (for the scaling factors of 2, 3.5 and 5).
@@
The time constant of the overall triplet population has been estimated by fitting scaled and shifted complimentary exponential decay curves to the purple data points in Figs. 14, 15 and 16, yielding $\tau_{2}^{\text{T}}=441.66$ fs, $\tau_{3.5}^{\text{T}}=28.44$ fs, and $\tau_{5}^{\text{T}}=12.44$ fs. Extrapolation of these time constants to $\alpha=1$ gives $\tau_{1}^{\text{T}}=6.06$ ps (as shown in Fig. 17). The inverse of $\tau_{1}^{\text{T}}$ is the ISC rate constant $k_{\text{ISC}}=1.73\times 10^{11}$ s-1.
@@
The lifetime of $\text{S}_{1}$ has been estimated by fitting scaled and shifted exponential decay curves to the yellow data points in Figs. 14, 15 and 16 (see also Figs. S25, S26 and S27), yielding $\tau_{2}^{\text{S}_{1}}=179.1$ fs, $\tau_{3.5}^{\text{S}_{1}}=26.3$ fs, and $\tau_{5}^{\text{S}_{1}}=8.8$ fs. Extrapolation of these lifetimes to $\alpha=1$ gives $\tau_{1}^{\text{S}_{1}}=1.73$ ps (as shown in Fig. S28).
@@
In order to explicitly show the importance of the included higher-lying excited states, the average population of the first three triplets has been separated from that of the higher-lying triplets (T4 through T10) in Figs. 18, 19 and 20, one for each value of the scaling factor $\alpha$. From the figures it can be seen that the higher-lying triplets are becoming strongly populated on the same time scale as the lower-lying ones, while the higher-lying singlets gain a non-negligible population as well.
Figure 18: Average populations of singlets and triplets (back-transformed from the spin-adiabatic basis) in the excited-state dynamics of I-BODIPY started in S1 for the scaling factor $\alpha=2$.
Figure 19: Average populations of singlets and triplets (back-transformed from the spin-adiabatic basis) in the excited-state dynamics of I-BODIPY started in S1 for the scaling factor $\alpha=3.5$.
Figure 20: Average populations of singlets and triplets (back-transformed from the spin-adiabatic basis) in the excited-state dynamics of I-BODIPY started in S1 for the scaling factor $\alpha=5$.
The large populations of the higher-lying singlets and triplets may seem surprising and can even raise suspicions that the law of conservation of energy has been violated. Actually, the average kinetic energy of about 3.2 eV distributed among the atomic nuclei of I-BODIPY at the beginning of each trajectory, the energy of an absorbed photon, and also some part of the average potential energy of the molecule in its ground state (in the initial geometry) formally transferred into the first excited singlet is more than enough to power nonadiabatic transitions to all the pure singlets and triplets included in our multielectron basis, at least at the employed TD-DFT(TDA)/BHLYP level of theory, cf. Tables S7a–c. Moreover, these and (hypothetically) even higher-lying spin-diabatic (i.e., scalar relativistic) states can be partially populated through their admixtures in some lower-lying spin-adiabatic (i.e., quasirelativistic) states. However, from Tables S2a–c it follows that if we switch from TD-DFT to ADC(2), the energy of S1 with respect to S0 decreases as a result of which all the vertical energy gaps between S1 and the higher-lying states increase. That is why we think that the average populations of the higher-lying spin-diabatic states observed in our TSH MD simulations are likely to be somewhat overestimated.
Eventually, all the simulations have been carried out on a single (nonrotating) molecule (whose center of mass is at rest) put in an empty space. In reality, however, the molecule interacts with the environment, which leads, though probably on a longer timescale than is the length of our trajectories, to a dissipation of its vibrational energy into its rotational and translational degrees of freedom as well as into all forms of energy of the surrounding solvent molecules not included in our model. As a result, the average singlet and triplet populations in the real molecule are assumed to continuously flow from the higher-lying to the lower-lying electronic states “after we stop watching”, supposedly on a longer than picosecond timescale.
@@
However, it should be noted that the data shown in Figs. S38 and 22 are not completely accurate or even correct. The problem is that by picking out only the time steps at which surface hops between spin-adiabatic states take place we obviously miss the transitions between spin-diabatic states which occur by continuous change of their CI expansion coefficients in the (one and only) active spin-adiabatic state. Actually, sufficiently close to a conical intersection of two spin-adiabatic states (crossing seam of two spin-diabatic states) a hop between spin-adiabatic states may in fact represent staying in a given spin-diabatic state (either singlet or triplet) and vice versa, staying in a given spin-adiabatic state may represent a hop between spin-diabatic states (from singlet to triplet or the other way around). Another undesired consequence of our slightly biased approach are the non-empty histograms on the diagonal of the matrix on Fig. S38.